OpenStudy (anonymous):

how to solve x^ + 1x - 72 = 0 ?? using quadratic formula or factoring ???

OpenStudy (anonymous):

\[x^2+x-72=0\] \[(x-8)(x+9)=0\] \[x-8=0\] \[x=8\] or \[x+9=0\] \[x=-9\]

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

welcome